Barnsley District Command Team


In Attendance:             Chief Superintendent Liz Watson (Chair)

Superintendent Shelley Hemsley, Operations

Mary Verity, Business Manager

Lisa Ingham, OMA

Amanda Tankard, HR Operational Partner


Apologies:                   CI Andy Hodgkinson, Response

DCI Debbie Mahmood, Crime Manager


1.         APOLOGIES


  1. As above.





2.1       The minutes of the last meeting dated 12 May 2015 were agreed as an accurate record.





3.1        Closed.


3.2        CrimingDCI Mahmood to brief PCT on the importance of criming immediates/from the scene.  Completed.  Discharged.



4.         LPT MOVE


4.1        Lockers and desks are currently being transferred, with checks on the fire alarm panel taking place.  A locker resource meeting will take place today; teams will be contacted in respect of the issue of locker keys and Airwave. 


4.2        The Diamond Team will attend District on 4th June 2015 for a final meet prior to the go live date.


4.3        GRS has been updated to reflect the LPU moves.  The siting of specialist role officers within GRS was identified within today’s meeting.  Following the VER outcome, PCSOs will be re-aligned within District.  Supt. Hemsley advised she would be reviewing flexible working arrangements for the PCSOs.


4.4        Command support over the weekend prior to the move and Command Team members attending briefings will be determined following this meeting. 


4.5        A site visit will take place at Churchfield Court in respect of security cameras and police signage.


4.6        Work at Wombwell will be completed by Friday of this week.  IS will relocate all the computers on Monday 1st June and at the same time install the briefing Logitec camera.



5.         VER UPDATE


5.1        District Commanders will receive a full list of authorised VER applicants on Tuesday 2nd June in order to notify staff.  The extending of VER officers leaving SYP employ by 30th June will be further discussed.





6.1        Summary of Year to Date Performance re. PCC Targets


The average monthly volume figures were reviewed.  Looking at the CSP comparison for the most similar groups, Barnsley is performing very well and is within the averages of that group.  In respect of violence against the person, although there is an increase across the Force and at Barnsley, District are still below the numbers in the most similar groups.  The only other exception is all other theft which covers makes offs, theft from gardens etc. which is just on the higher limit of the averages, i.e. 0.8 crimes per 1k population.


In terms of the recent summary of performance over the last month, during April there was a reduction in burglaries.  Violence in public has increased, over the last 90 days District had a spike on the weekends covering the 16th May and 25th May (BH weekend), with an average of 4-10 offences on these dates.  Nationally there has been in an increase in public violence. 


6.2        Forthcoming Events/Operations


The LPT restructure will take place on 8th June 2015.


The High Sheriff will patrol with Special Constables the weekend of the 13th June.


An afternoon music festival will take place on the 20th June 2015.





7.1        Major Crime Investigations


Deferred to the next meeting.





8.1        Update from Key Force Meetings


            JCB.  There will be a joint collaboration board within JCB with the two Chief Constables and Crime Commissioners attending.  Baselining will take place of different areas.  The Board will guideline the financial floor, i.e. identify minimum savings with an expectation that departments will  plan further savings above this figure.  Subject matter experts will be the lead review officers with a core team of support.  Benchmarking is expected to take around 6 months with corporate teams currently being advertised. 


            LPIB.  ACC Byrne discussed ongoing Force-wide work around violent crime.  Binge will now be named Op Tandem with the night shift taking over responsibility from the afternoon team.  Work will take place around violent crime in terms of contact with the victim.  Officers should try and contact the victim, however if there is no response then a Legal Services approved letter should be left at the victim’s home address outlining the crime number with advice that the incident will be filed unless the person makes further contact with SYP.


            Officers will be able to file incidents a lot quicker if there are no further lines of enquiry, which CRB can progress. 


            The floor walker officers will remain at Alas Court longer than the anticipated end of May date; a phased return will be negotiated with Districts.  The officers have been requested to brief the Command Team and Inspectors meeting with their findings around reducing demand.


            The Doncaster LPU restructure is in place with only minor issues identified. 




8.2        Closed.


8.3        Closed.


8.4        Business Plan


Business Plan Projects.  All have been updated. 


  • Accommodation strategy development.  Mary Verity is currently reviewing and updating the accommodation plan which feeds into the Force strategy.  Co-location opportunities were discussed.
  • Potential co-location of BMBC.  A preferred option has been confirmed in terms of the layout of the accommodation.  FM are identifying costs from contractors with the work to take place between June and October 2015.
  • Continuous improvement.  A meeting will take place on 18th June 2015.
  • Vulnerable Witness Suite at Dodworth.  Progression of the work is delayed. 


8.5       Human Resource Issues




Police officers – permanent 257, FTE 252.53 with 18 temporaries (17.95 FTE).  Seven are externally funded.  Strength 262, FTE 255.71.  Inspector Batty to link in with Amanda Tankard in respect of current numbers now that GRS has been updated.


Police staff – 32, head count 27.39 FTE with 1 temporary post.


Personnel & Welfare Issues.  Discussed.




8.6        Performance & Governance Meeting Action Schedule


Deferred to the next meeting.


8.7        Audit Reports


            There are no outstanding audit reports.


8.8        Closed.





7.1       None.


Tuesday, 26 May, 2015 - 10:00